# New Foundations is consistent

### v22

@@
New Foundations is consistent
M. Randall Holmes and Sky Wilshaw
@@ 1.1 Introductory note
The first author showed the equiconsistency of NF with the quite bizarre system TTT (Tangled Type Theory) in 1995 [see [6], but this paper is not really recommended; the presentation of its results here is better], which gave a possible approach to a consistency proof. Since 2010, the first author has been attempting to write out proofs, first of the existence of a tangled web of cardinals in Mac Lane set theory [an indication of this approach is given below] and then directly of the existence of a model of tangled type theory. These proofs are difficult to read, insanely involved, and involve the sort of elaborate bookkeeping which makes it easy to introduce errors every time a new draft is prepared. The second author (with the assistance of others initially) has formally verified the proof of the first author of the existence of a model of TTT (see [24]), and so of the consistency of New Foundations, in the Lean proof verification system, and in the process has suggested minor corrections and considerable formal improvements to the argument originally proposed by the first author. The second author reports that the formalized proof is still difficult to read and insanely involved with nasty bookkeeping. Both authors feel that there ought to be a simpler approach, but the existing argument at least strongly resists attempts at simplification.
@@
Any remarks in the first person singular may be attributed to the first author.
@@ 1.2 Acknowledgements
1.3 Dated remarks on versions
initial note for this series: The formal verification of the results in section 3 and 4 is complete. Sky’s Lean formalization concludes with a proof that there is a structure for the language of TTT which satisfies all typed versions of the statements in the Hailperin axiomatization of NF. Such a structure is a model of TTT, and the existence of a model of TTT implies the consistency of New Foundations. Thus we know that the construction in the paper is correct in general terms and we need to rewrite it for publication to parallel the formalized proof as much as possible: we certainly know that the conclusion of the paper expressed in the title is true.
initial note for the previous series: This document is probably my best overall version so far. The immediate occasion for its preparation was to serve students attempting to verify the proof in Lean. A formal verification should in our view (at least in the view of the first author) avoid metamathematics as far as possible, so it is the fact that the structure defined in section 3 is a model of TTT which should be verified, and further, a finite axiomatization (mod type indexing) of TST and thus TTT should be verified in the model in lieu of the usual statement of the axiom of comprehension of TTT. This program has been completed, on the Lean side.
9/30/2024:
This version will be posted to arxiv today. Changes from 9/10 are minor. A lot of the version notes have been edited out, and all the editorial remarks, as arxiv requested.
Mod the possibility of further remarks from Sky, this is the new version for Arxiv. Arxiv readers should note that the proof of Freedom of Action is completely rewritten by Sky Wilshaw in this version. It is in some sense the same argument as my original argument, or close to being so, but it is much easier to read. The last of the series of drafts before Sky’s revision of the FoA proof is preserved on my web page for the time being.
9/3/2024:
Starting new version for Arxiv with editing comments suppressed, reflecting the rewrite of the proof of Freedom of Action by the second author and various minor corrections and rearrangements.
8/16/2024:
New revisions by Sky and Holmes responses. Extended type indices now always have minimum $-1$.
8/13/2024:
Working copy for the meeting with Sky on this date.
8/11/2024:
This version incorporates a substantial bookkeeping change. The second components of support conditions now have $-1$ as minimal element, which improves uniformity in some boundary cases, but which necessitates a lot of changes to indexing and notation here and there.
7/26/2024:
This contains Sky’s complete rewrite of the proof of Freedom of Action, without changes to subsequent text which it entails. It also includes redefinition of the position of function to act on singletons of atoms instead of atoms, to avoid irritating type conflicts between the definition of the position function and the definition of support conditions.
7/2/2024:
Starting a revision effort for the month of July, to support reconciliation of the paper with the formal proof with an eye to submission for publlcation.
@@ 2.1 The simple theory of types TST and TSTU
We introduce a theory which we call the simple typed theory of sets or TST, a name favored by the school of Belgian logicians who studied NF (théorie simple des types). This is not the same as the simple type theory of Ramsey and it is most certainly not Russell’s type theory (see historical remarks below).
@@
allowing many objects with no elements (called atoms or urelements) in each positive type. A technically useful refinement adds a constant $\emptyset^{i}$ of each positive type $i$ with no elements: we can then address the problem that $\{x^{i}:\phi\}$ is not necessarily uniquely defined when $\phi$ is uniformly false by defining $\{x^{i}:\phi\}$ as $\emptyset^{i+1}$ in this case.
@@ 2.1.1 Typical ambiguity
TST(U) exhibits a symmetry which is important in the sequel.
@@ 2.1.2 Historical remarks
In 1914 ([23]), Norbert Wiener gave a definition of the ordered pair as a set (not the one now in use) and seems to have recognized that the type theory of Principia Mathematica could be simplified to something like TST, but he did not give a formal description. The theory we call TST was apparently first described by Tarski in the 1930s.
@@
2.2 Some mathematics in TST and some subtheories of interest
@@
2.2.1 The theories TSTn and their natural models
For each natural number $n$, the theory TSTn is defined as the subtheory of TST with vocabulary restricted to use variables only of types less than $n$ (TST with $n$ types).
We note the existence of models of TST and of each TSTn in ordinary set theory in which the set $X_{i+1}$ implementing type $i+1$ is the power set of the set $X_{i}$ implementing type $i$ and the membership of type $i$ objects in type $i+1$ objects is implemented by the restriction of the membership relation of the metatheory to $X_{i}\times X_{i+1}$. We call a model of TST or TSTn natural iff it is isomorphic to such a model. The exact models we describe have the inconvenient feature that the sets implementing the types have some overlap. Bounded Zermelo set theory implies the existence of natural models of TSTn for each concrete $n$; a not terribly strong extension of Zermelo set theory implies the existence of natural models of TST.
Further, each TSTn has what we may call natural models in TST itself, though some care must be exercised in defining them. Let $X$ be a set. Implement type $i$ for each $i<n$ as $\iota^{(n-1)-i}``{\cal P}^{i}(X)$. If $X$ is in type $j$, each of the types of this interpretation of TSTn is a set in the same type $j+n-1$. For any relation $R$, define $R^{\iota}$ as {($\{(\{x\},\{y\}):xRy\}$. The membership relation of type $i-1$ in type $i$ in the interpretation described is the restriction of $\subseteq^{\iota^{(n-1)-i}}$ to the product of the sets implementing type $i-1$ and type $i$.
Notice then that, for each concrete natural number $n$, we can define truth for formulas in these natural models of TSTn in TST, though not in a uniform way which would allow us to define truth for formulas in TST in TST.
Further, both in ordinary set theory and in TST, observe that truth of sentences in natural models of TSTn is completely determined by the cardinality of the set used as type 0, since two natural models of TST or TSTn with base types implemented by sets of the same cardinality are clearly isomorphic.
@@ 2.3 New Foundations and NFU
Quine’s observation is that since any theorem $\phi$ of TST is accompanied by theorems $\phi^{+},\phi^{++},\phi^{+++},\ldots$ and every defined object $\{x:\phi\}$ is accompanied by $\{x^{+}:\phi^{+}\},\{x^{++}:\phi^{++}\},\{x^{+++}:\phi^{+++}\}$, so the picture of what we can prove and construct in TST looks rather like a hall of mirrors, we might reasonably (?) suppose that the types are all the same.
@@
In the original paper, this is expressed in a way which avoids explicit dependence on the language of another theory. Let $\phi$ be a formula of the language of NF. A function $\sigma$ is a stratification of $\phi$ if it is a (possibly partial) map from variables to non-negative integers such that for each atomic subformula ‘$x=y$’ of $\phi$ we have $\sigma($$x$$)=\sigma($$y$)))) and for each atomic subformula ‘$x\in y$’ of $\phi$ we have $\sigma($$x$$)+1=\sigma($$y$)))). A formula $\phi$ is said to be stratified iff there is a stratification of $\phi$. Then for each stratified formula $\phi$ of the language of NF and variable $x$ we have an axiom $(\exists A:(\forall x:x\in A\leftrightarrow\phi))$. The stratified formulas are exactly the formulas $\phi^{*}$ up to renaming of variables.
NF has been dismissed as a “syntactical trick” because of the way it is defined. It might go some way toward dispelling this impression to note that the stratified comprehension scheme is equivalent to a finite collection of its instances, so the theory can be presented in a way which makes no reference to types at all. This is a result of Hailperin ([4]), refined by others, and discussed in some detail in a subsequent subsection of this section, because it is relevant to the Lean formalization. One obtains a finite axiomatization of NF by analogy with the method of finitely axiomatizing von Neumann-Gödel-Bernays predicative class theory. It should further be noted that the first thing one does with any finite axiomatization is prove stratified comprehension as a meta-theorem, in practice, but it remains significant that the theory can be axiomatized with no reference to types at all. It is also worth noting that the collection of all typed instances of the Hailperin axioms is an axiomatization of TST, not a finite axiomatization, but an axiomatization with finitely many templates.
@@
(uniqueness follows by extensionality) whch we denote by $\{x:\phi\}$.
@@
Jensen showed that NFU is consistent and moreoever NFU + Infinity + Choice is consistent. We will give an argument similar in spirit though not the same in detail for the consistency of NFU in the next section.
@@ 2.4 Tangled type theory TTT and TTTU
For any formula $\phi$ of the language of TST and any strictly increasing sequence $\{s_{i}\}_{\mathbb{N}}$ in $\lambda$, let $\phi^{s}$ be the formula obtained by replacing each variable of type $i$ with a variable of type $s(i)$. To make this work rigorously, we suppose that we have a bijection from type $i$ variables of the language of TST to type $\alpha$ variables of the language of TTT for each natural number $i$ and ordinal $\alpha<\lambda$.
TTT is then the first order theory with types indexed by the ordinals below $\lambda$ whose well-formed atomic sentences ‘$x=y$’ have ${\tt type}($$x$$)={\tt type}($$y$)))) and whose atomic sentences ‘$x\in y$’ satisfy ${\tt type}($$x$$)<{\tt type}($$y$)))), and whose axioms are the sentences $\phi^{s}$ for each axiom $\phi$ of TST and each strictly increasing sequence $s$ in $\lambda$. TTTU has the same relation to TSTU (with the addition of constants $\emptyset^{\alpha,\beta}$ for each $\alpha<\beta<\lambda$ such that $(\forall x^{\alpha}:x^{\alpha}\not\in\emptyset^{\alpha,\beta})$ is an axiom).
@@
    5. (e)
      But this implies by compactness that the full Ambiguity Scheme $\phi\leftrightarrow\phi^{+}$ is consistent with TST(U). We could at this point say that NF(U) is consistent by the 1962 result of Specker ([20]), but we will continue and indicate how to prove this directly.
    6. (f)
      If we have a model of TST(U) augmented with a Hilbert symbol [a primitive term construction $(\epsilon x:\phi)$ (same type as $x$) with axiom scheme $\phi[(\epsilon x:\phi)/x]\leftrightarrow(\exists x:\phi)$ which cannot appear in instances of comprehension (the quantifiers are not defined in terms of the Hilbert symbol, because they do need to appear in instances of comprehension)] and Ambiguity (for all formulas, including those which mention the Hilbert symbol) then we can readily get a model of NF(U), by constructing a term model using the Hilbert symbol in the natural way, then identifying all terms with their type-raised versions (an element of the term model is a class of terms $(\epsilon x:\phi)$ in the language of TST(U) which is obtained from a single such term by all possible upward or downward shifts of type by a constant amount: we write this $[(\epsilon x:\phi)]$).
    7. (g)
      All statements in the resulting type-free theory can be decided: the truth value of a closed atomic formula $[(\epsilon x:\phi)]\,R\,[(\epsilon y:\psi)]$ in the model of NF(U) is determined by replacing each side of the formula with one of its elements in such a way as to get a well-typed formula, which will always be possible if types are taken high enough, and reading the truth value of the resulting formula from the model of TST(U) with Ambiguity and Hilbert symbol; $R$ is either = or $\in$ [note that the relation interpreting equality in the term model is a nontrival equivalence relation]. Truth values of more complex statements are determined in standard ways. That this makes axioms of NF(U) true should be clear.
    8. (h)
      Now observe that a model of TTT(U) can readily be equipped with a Hilbert symbol if this creates no obligation to add instances of comprehension containing the Hilbert symbol (use a well-ordering of the set implementing each type to interpret a Hilbert symbol $(\epsilon x:\phi)$ in that type as the first $x$ such that $\phi$), and the argument in paragraphs (a)–(e) adapts to show consistency of TST(U) plus Ambiguity with the Hilbert symbol.
    9. (i)
@@
    10. (j)
      From this it follows that NF(U) is consistent.
@@
It is enough to exhibit a model of TTTU. Suppose $\lambda>\omega$. Represent type $\alpha$ as $V_{\omega+\alpha}\times\{\alpha\}$ for each $\alpha<\lambda$ ($V_{\omega+\alpha}$ being a rank of the usual cumulative hierarchy). Define $\in_{\alpha,\beta}$ for $\alpha<\beta<\lambda$ as
@@
This gives a model of TTTU in which the membership of type $\alpha$ in type $\beta$ interprets each $(y,\beta)$ with $y\in V_{\omega+\beta}\setminus V_{\omega+\alpha+1}$ as an urelement.
Our use of $V_{\omega+\alpha}$ enforces Infinity in the resulting models of NFU (note that we did not have to do this: if we set $\lambda=\omega$ and interpret type $\alpha$ using $V_{\alpha}$ we prove the consistency of NFU with the negation of Infinity). It should be clear that Choice holds in the models of NFU eventually obtained if it holds in the ambient set theory.
@@
Of course Jensen did not phrase his argument in terms of tangled type theory. Our contribution here was to reverse engineer from Jensen’s original argument for the consistency of NFU an argument for the consistency of NF itself, which requires additional input which we did not know how to supply (a proof of the consistency of TTT itself). An intuitive way to say what is happening here is that Jensen noticed that it is possible to skip types in a certain sense in TSTU in a way which is not obviously possible in TST itself; to suppose that TTT might be consistent is to suppose that such type skipping is also possible in TST.
2.4.1 How internal type representations unfold in TTT
We have seen above that TST can internally represent TSTn (on page 2.2.1). An attempt to represent types of TTT internally to TTT has stranger results. The development of the model does not depend on reading this section.
In TST the strategy for representing type $i$ in type $n\geq i$ is to use the $n-i$-iterated singleton of any type $i$ object $x$ to represent $x$; then membership of representations of type $i-1$ objects in type $i$ objects is represented by the relation on $n-i$-iterated singletons induced by the subset relation and with domain restricted to $n-(i+1)$-fold singletons. This is described more formally above.
In TTT the complication is that there are numerous ways to embed type $\alpha$ into type $\beta$ for $\alpha<\beta$ along the lines just suggested. We define a generalized iterated singleton operation: where $A$ is a finite subset of $\lambda$, $\iota_{A}$ is an operation defined on objects of type ${\tt min}(A)$. $\iota_{\{\alpha\}}(x)=x$. If $A$ has $\alpha<\beta$ as its two smallest elements, $\iota_{A}(x)$ is $\iota_{A_{1}}(\iota_{\alpha,\beta}(x))$, where $A_{1}$ is defined as $A\setminus\{{\tt min}(A)\}$ (a notation we will continue to use) and $\iota_{\alpha,\beta}(x)$ is the unique type $\beta$ object whose only type $\alpha$ element is $x$.
Now for any nonempty finite $A\subseteq\lambda$ with minimum $\alpha$ and maximum $\beta$, the range of $\iota_{A}$ is a set, and a representation of type $\alpha$ in type $\beta$. For simplicity we carry out further analysis in types $\beta,\beta+1,\beta+2\ldots$ though it could be done in more general increasing sequences. Use the notation $\tau_{A}$ for the range of $\iota_{A}$, for each set $A$ with $\beta$ as its maximum. Each such set has a cardinal $|\tau_{A}|$ in type $\beta+2$. It is a straightforward argument in the version of TST with types taken from $A$ and a small finite number of types $\beta+i$ that $2^{|\tau_{A}|}=|\tau_{A_{1}}|$ for each $A$ with at least two elements. The relevant theorem in TST is that $2^{|\iota^{n+1}``X|}=\iota^{n}``X$, relabelled with suitable types from $\lambda$. We use the notation $\exp(\kappa)$ for $2^{\kappa}$ to support iteration. Notice that for any $\tau_{A}$ we have $\exp^{|A|-1}(|\tau_{A}|)=|\tau_{\{\beta\}}|$, the cardinality of type $\beta$. Now if $A$ and $A^{\prime}$ have the same minimum $\alpha$ and maximum $\beta$ but are of different sizes, we see that $|\tau_{A}|\neq|\tau_{A^{\prime}}|$, since one has its $|A|-1$-iterated exponential equal to $|\tau_{\{\beta\}}|$ and the other has its $|A^{\prime}|-1$-iterated exponential equal to $|\tau_{\{\beta\}}|$. This is odd because there is an obvious external bijection between the sets $\tau_{A}$ and $\tau_{A^{\prime}}$: we see that this external bijection cannot be realized as a set. $\tau_{A}$ and $\tau_{A^{\prime}}$ are representations of the same type, but this is not obvious from inside TTT. We recall that we denote $A\setminus\{{\tt min}(A)\}$ by $A_{1}$; we further denote $(A_{i})_{1}$ as $A_{i+1}$. Now suppose that $A$ and $B$ both have maximum $\beta$ and $A\setminus A_{i}=B\setminus B_{i}$, where $i<|A|\leq|B|$. We observe that for any concrete sentence $\phi$ in the language of TSTi, the truth value of $\phi$ in natural models with base type of sizes $|\tau_{A}|$ and $|\tau_{B}|$ will be the same, because the truth values we read off are the truth values in the model of TTT of versions of $\phi$ in exactly the same types of the model (truth values of $\phi^{s}$ for any $s$ having $A\setminus A_{i}=B\setminus B_{i}$ as the range of an initial segment). This much information telling us that $\tau_{A_{j}}$ and $\tau_{B_{j}}$ for $j<i$ are representations of the same type is visible to us internally, though the external isomorphism is not. We can conclude that the full first order theories of natural models of TSTi with base types $|\tau_{A}|$ and $|\tau_{B}|$ are the same as seen inside the model of TTT, if we assume that the natural numbers of our model of TTT are standard (which is quite a strong assumption!).
2.4.2 Tangled webs of cardinals: a suggestion of another approach not followed here
Nothing in the construction of a model of tangled type theory and verification that it is a model which appears below depends on anything in this section.
It is straightforward to transform a model of TST into a model of bounded Zermelo set theory (Mac Lane set theory) with atoms or without foundation (this depends on how type 0 is handled). Specify an interpretation of type 0 either as a set of atoms or a set of self-singletons. Then interpret type $i+1$ as inhabited by sets of type $i$ objects in the obvious way, identifying type $i+1$ objects with objects of lower type which happen to have been assigned the same extension.
In a model of TTT, do this along some increasing sequence of types of order type $\omega$ whose range includes an infinite ordinal $\alpha$. In the resulting model of bounded Zermelo set theory, let $\tau_{A}$ represent the cardinality of the range of $\iota_{A}$ as in the previous discussion (for nonempty subsets of type $A$ all with maximum the same infinite $\alpha$). Suppose further for the sake of argument that our model of TTT is $\lambda$-complete, in the sense that any subset of a type of cardinality that of $\lambda$ or less is implemented as a set in each higher type. It will follow that $A\mapsto\tau_{A}$ is actually a function. [It is an incidentally interesting fact that the models we construct (with no dependence on this section) can have this completeness property (using section 3 terminology, under the condition $\lambda<\kappa$ we have $\lambda$-completeness; $\alpha<\kappa$ actually suffices for existence of a tangled web)].
We describe the situation which holds for these cardinals.
We work in Mac Lane set theory. Choice is not assumed, and we use the Scott definition of cardinals.
Definition 2.3.
If $A$ is a nonempty finite set of ordinals which is sufficiently large, we define $A_{1}$ as $A\setminus{\tt min}(A)$ and $A_{0}$ as $A$, $A_{i+1}$ as $(A_{i})_{1}$.
Definition 2.4.
A tangled web of cardinals of order $\alpha$ ($\alpha$ an infinite ordinal) is a function $\tau$ from the set of nonempty sets of ordinals with $\alpha$ as maximum to cardinals such that
  1. 1.
    If $|A|>1$, $\tau(A_{1})=2^{\tau(A)}$.
  2. 2.
    If $|A|\geq n$, the first order theory of a natural model of TSTn with base type $\tau(A)$ is completely determined by $A\setminus A_{n}$, the $n$ smallest elements of $A$.
The bookkeeping in different versions of this definition in different attempts at a tangled web version of the proof of the consistency of NF have been different (an obvious point about the version given here is that the top ordinal $\alpha$ could be omitted). Another remark is that it is clear that asserting the existence of a tangled web is stronger than simple TTT: it requires $\lambda>\omega$, and the $\lambda$-completeness of course is a strong assumption in the background. All variants that I have used support versions of the following

### v23

@@
Abstract
In this paper we present a proof that Quine’s set theory New Foundations of [12], 1937 is consistent. We do this by constructing a model of the “tangled type theory” proposed by the first author and shown to be equiconsistent with NF in [6]. It is worth noting that the fact that the structure we describe is a model of tangled type theory has been formally verified in Lean by the second author and others (in [24]); this paper alludes occasionally to the Lean formalization (and relevant sections are parallel to it in structure), but the paper version does not depend on the formal verification.
@@ 1.1 Introductory note
The first author showed the equiconsistency of NF with the quite bizarre system TTT (Tangled Type Theory) in 1995 [see [6], but this paper is not really recommended; the presentation of its results here is better], which gave a possible approach to a consistency proof. Since 2010, the first author has been attempting to write out proofs, first of the existence of a “tangled web of cardinals” in Mac Lane set theory [not defined in this paper] and then directly of the existence of a model of tangled type theory. These proofs have been difficult to read, insanely involved, and involve the sort of elaborate bookkeeping which makes it easy to introduce errors every time a new draft is prepared. We hope that the present version is more pleasant, though we know it is still forbidding. The second author (with the assistance of others initially) has formally verified the proof of the first author of the existence of a model of TTT (see [24]), and so of the consistency of New Foundations, in the Lean proof verification system, and in the process has suggested minor corrections and considerable formal improvements to the argument originally proposed by the first author. The second author reports that the formalized proof is still difficult to read and insanely involved with nasty bookkeeping. Both authors feel that there ought to be a simpler approach, but the existing argument at least strongly resists attempts at simplification.
@@
Any remarks in the paper in the first person singular may be attributed to the first author.
@@ 2.1 The simple theory of types TST and TSTU
We introduce a theory which we call the simple typed theory of sets or TST, a name favored by the school of Belgian logicians111led by Maurice Boffa, to whom the first author owes a profound debt, for being harbored in his group at the beginning of my career, which made it possible for me to launch my research program without distractions attendant on looking for a job. who studied NF (théorie simple des types). This is not the same as the simple type theory of Ramsey and it is most certainly not Russell’s type theory (see historical remarks below).
@@
allowing many objects with no elements (called atoms or urelements) in each positive type. A technically useful refinement adds a constant $\emptyset^{i}$ of each positive type $i$ with no elements: we can then address the problem that $\{x^{i}:\phi\}$ is not necessarily uniquely defined when $\phi$ is uniformly false by defining $\{x^{i}:\phi\}$ as $\emptyset^{i+1}$ in this case. The superscript is sometimes omitted from $\emptyset^{i}$ (and other constants parametrized by types) when the type can be deduced from context.
@@ 2.1.1 Typical ambiguity
TST(U) exhibits a symmetry, often called typical ambiguity, which is important in the sequel.
@@ 2.1.2 Historical remarks
In 1914 ([23]), Norbert Wiener gave a definition of the ordered pair as a set (not the one now in use) and seems to have recognized that the type theory of Principia Mathematica could be simplified to something like TST, but he did not give a formal description. The theory we call TST was apparently first described by Tarski in the 1930s [^3] We regard it as an important feature of TST that the nature of type 0 is left completely undetermined, so we do not regard descriptions of arithmetic of order $\omega$ appearing about the same time as part of this history. Arithmetic of order $\omega$ is a similar typed theory but lacks the ambiguity of interest here..
@@
2.1.3 Some mathematics in TST
@@ 2.2 New Foundations and NFU
Quine’s observation is that since any theorem $\phi$ of TST is accompanied by theorems $\phi^{+},\phi^{++},\phi^{+++},\ldots$ and every defined object $\{x:\phi\}$ is accompanied by $\{x^{+}:\phi^{+}\},\{x^{++}:\phi^{++}\},\{x^{+++}:\phi^{+++}\}$, so the picture of what we can prove and construct in TST looks rather like a hall of mirrors [^5] I have been asked whether the turn of phrase “hall of mirrors” here is my own or due to Quine himself or others in the interim: I actually do not know!, we might reasonably (?) suppose that the types are all the same.
@@
In the original paper, this is expressed in a way which avoids explicit dependence on the language of another theory. We do this in a way which is usual now but not necessarily identical to what is done in the original paper. Let $\phi$ be a formula of the language of NF. A function $\sigma$ is a stratification of $\phi$ if it is a (possibly partial) map from variables to non-negative integers such that for each atomic subformula ‘$x=y$’ of $\phi$ we have $\sigma($$x$$)=\sigma($$y$)))) and for each atomic subformula ‘$x\in y$’ of $\phi$ we have $\sigma($$x$$)+1=\sigma($$y$)))). A formula $\phi$ is said to be stratified iff there is a stratification of $\phi$. Then for each stratified formula $\phi$ of the language of NF and variable $x$ we have an axiom $(\exists A:(\forall x:x\in A\leftrightarrow\phi))$. The stratified formulas are exactly the formulas $\phi^{*}$ up to renaming of variables.
@@
(uniqueness follows by extensionality) which we denote by $\{x:\phi\}$.
NF has been dismissed as a “syntactical trick” because of the way it is defined. It might go some way toward dispelling this impression to note that the stratified comprehension scheme is equivalent to a finite collection of its instances, so the theory can be presented in a way which makes no reference to types at all. This is a result of Hailperin ([4]), refined by others, and discussed in some detail in a subsequent subsection of this section, because it is relevant to the Lean formalization. One obtains a finite axiomatization of NF by analogy with the method of finitely axiomatizing von Neumann–Gödel–Bernays predicative class theory. It should further be noted that the first thing one does with any finite axiomatization is prove stratified comprehension as a meta-theorem, in practice, but it remains significant that the theory can be axiomatized with no reference to types at all. It is also worth noting that the collection of all typed instances of the Hailperin axioms is an axiomatization of TST, not a finite axiomatization, but an axiomatization with finitely many templates.
@@
Jensen showed that NFU is consistent and moreover NFU + Infinity + Choice is consistent. We will give an argument similar in spirit though not the same in detail for the consistency of NFU in the next section.
@@ 2.3 Tangled type theory TTT and TTTU
For any formula $\phi$ of the language of TST and any strictly increasing sequence $\{s_{i}\}_{i\in\mathbb{N}}$ in $\lambda$, let $\phi^{s}$ be the formula obtained by replacing each variable of type $i$ with a variable of type $s(i)$. To make this work rigorously, we suppose that we have a bijection from type $i$ variables of the language of TST to type $\alpha$ variables of the language of TTT for each natural number $i$ and ordinal $\alpha<\lambda$.
TTT is then the first order theory with types indexed by the ordinals below $\lambda$ whose well-formed atomic sentences ‘$x=y$’ have ${\tt type}($$x$$)={\tt type}($$y$)))) and whose well-formed atomic sentences ‘$x\in y$’ satisfy ${\tt type}($$x$$)<{\tt type}($$y$)))), and whose axioms are the sentences $\phi^{s}$ for each axiom $\phi$ of TST and each strictly increasing sequence $s$ in $\lambda$. TTTU has the same relation to TSTU (with the addition of constants $\emptyset^{\alpha,\beta}$ for each $\alpha<\beta<\lambda$ such that $(\forall x^{\alpha}:x^{\alpha}\not\in\emptyset^{\alpha,\beta})$ is an axiom).
@@
  4. 4.
    But this implies by compactness that the full Ambiguity Scheme $\phi\leftrightarrow\phi^{+}$ is consistent with TST(U). We could at this point say that NF(U) is consistent by the 1962 result of Specker ([20]), but we will continue and indicate how to prove this directly (our second aim).
  5. 5.
    We pause to add a useful notion to our logic. The Hilbert symbol is a primitive term construction $(\epsilon x:\phi)$ (same type as $x$) with axiom schemes $(\exists x:\phi)\rightarrow\phi[(\epsilon x:\phi)/x]$ and $(\forall x:\phi\leftrightarrow\psi)\rightarrow(\epsilon x:\phi)=(\epsilon x:\psi)$ which cannot appear in instances of comprehension. We do not define the quantifiers in terms of the Hilbert symbol, because the quantifiers do need to appear in instances of comprehension.
  6. 6.
    If we have a model of TST(U) augmented with a Hilbert symbol which satisfies Ambiguity (for all formulas, including those which mention the Hilbert symbol) then we can readily get a model of NF(U), by constructing a term model using the Hilbert symbol in the natural way, then identifying all terms with their type-raised versions (an element of the term model is a class of terms $(\epsilon x:\phi)$ in the language of TST(U) which is obtained from a single such term by all possible upward or downward shifts of type by a constant amount: we write this $[(\epsilon x:\phi)]$).
  7. 7.
    All statements in the resulting type-free theory can be decided: the truth value of a closed atomic formula $[(\epsilon x:\phi)]\,R\,[(\epsilon y:\psi)]$ in the model of NF(U) is determined by replacing each side of the formula with one of its elements in such a way as to get a well-typed formula, which will always be possible if types are taken high enough, and reading the truth value of the resulting formula from the model of TST(U) with Ambiguity and Hilbert symbol; $R$ is either = or $\in$ [note that the relation interpreting equality in the term model is a nontrivial equivalence relation]. Truth values of more complex statements are determined in standard ways. That this makes axioms of NF(U) true should be clear.
  8. 8.
    Now observe that a model of TTT(U) can readily be equipped with a Hilbert symbol if this creates no obligation to add instances of comprehension containing the Hilbert symbol (use a well-ordering of the set implementing each type to interpret a Hilbert symbol $(\epsilon x:\phi)$ in that type as the first $x$ such that $\phi$ or as a default object of the appropriate type if there is no such $x$), and the argument in paragraphs (1)–(4) adapts to show consistency of TST(U) plus Ambiguity with the Hilbert symbol.
  9. 9.
@@
  10. 10.
    From this it follows that NF(U) is consistent under the hypothesis that TTT(U) is consistent, which is what was to be shown.
@@
It is enough to exhibit a model of TTTU. Represent type $\alpha$ as $V_{\omega+\alpha}\times\{\alpha\}$ for each $\alpha<\lambda$ ($V_{\omega+\alpha}$ being the rank indexed by $\omega+\alpha$ in the usual cumulative hierarchy). Define $\in_{\alpha,\beta}$ for $\alpha<\beta<\lambda$ as
@@
This gives a model of TTTU in which the membership of type $\alpha$ in type $\beta$ is defined in such a way that each $(y,\beta)$ with $y\in V_{\omega+\beta}\setminus V_{\omega+\alpha+1}$ is interpreted as an urelement over type $\alpha$.
Our use of $V_{\omega+\alpha}$ enforces Infinity in the resulting models of NFU (note that we did not have to do this: if we set $\lambda=\omega$ and interpret type $\alpha$ using $V_{\alpha}$ we prove the consistency of NFU with the negation of Infinity). It should be clear that Choice (stated in a stratified way) holds in the models of NFU eventually obtained if it holds in the ambient set theory.
@@
Of course Jensen did not phrase his argument in terms of tangled type theory. Our contribution here was to reverse engineer from Jensen’s original argument for the consistency of NFU an argument for the consistency of NF itself, which requires additional input which we did not know how to supply (a proof of the consistency of TTT itself). An intuitive way to say what is happening here is that Jensen noticed that it is possible to skip types in a certain sense in TSTU in a way which is not obviously possible in TST itself; to suppose that TTT might be consistent is to suppose that such type skipping is also possible in TST, and the description of TTT is an explicit realization of what is meant by this. We leave it to the historically inclined reader to examine Jensen’s original proof and determine the exact relationship between the approaches.
2.3.1 There are $\alpha$-models for every $\alpha$
This section is devoted to adapting a result of Jensen to our context. Jensen proved in [10] that for any infinite ordinal $\alpha$ of the metatheory, there is a model of NFU + Infinity in which there is a well-ordering of type $\alpha$. This is of particular interest in the case $\alpha=\omega$, which establishes that the natural numbers of a model can correspond precisely to the natural numbers of the metatheory.
Nothing in our construction of a model of TTT below depends on this section: the fact that an existence of an $\omega$-model of NF follows from our main result is significant for extensions of our results and uses this theorem.
A relation $R$ of one of the theories under discussion is a set of ordered pairs [^6] The discussion here presupposes use of the Kuratowski ordered pair; it could be adapted to the use of another pair definition. of the theory in question; note that it corresponds to a set of ordered pairs of the metatheory with the same field. If this relation corresponding to $R$ is a well-ordering, we can talk about its order type both in terms of the theory under consideration and in terms of the metatheory, though ordinals of the theory under consideration and the metatheory may not be the same.
In any of the theories under discussion, an ordinal is taken to be an equivalence class of well-orderings under similarity. The relations which belong to an ordinal of one of these theories may fail to be well-orderings in terms of the metatheory. If $\alpha$ is an ordinal of the metatheory, we say that $\alpha$ is implemented in a type of one of the theories under discussion exactly if there is an ordinal in that type (in the sense of the theory) whose elements in fact correspond to well-orderings of order type $\alpha$ in the sense of the metatheory, in the sense outlined in the previous paragraph. It is straightforward to show that if $\alpha$ is implemented in a type, the ordinal implementing it will contain all well-orderings of appropriate type in the sense of the theory whose order type is $\alpha$ in the sense of the metatheory.